OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). sin 2x - sin 4x = 0

OpenStudy (anonymous):

sin 2x = sin 4x \[\sin (n\pi-x)=\sin(x) : n \in {Z} \]